OpenStudy (anonymous):

How does -5(2m-3)-(y-2) =-11y+17? Is this a type-o or am I missing something?

OpenStudy (anonymous):

I don't see anything wrong with it...

OpenStudy (anonymous):

There is an infinite number of y and m value pairs that would satisfy this equation, and you only need one to show that it is true. For example, you could use m=5 and y=5.

OpenStudy (anonymous):

I'm new at this but I came up with -10m+17-y

OpenStudy (anonymous):

No you did something wrong. Does your problem have instructions? What are you trying to solve for?

OpenStudy (anonymous):

Oh so in this case could you show me the steps taken

OpenStudy (anonymous):

simplify any expression that can be simplified

OpenStudy (anonymous):

were the instructions

OpenStudy (anonymous):

Oh okay. So for this one, these are the steps you could take:\[-5(2m-3)-(y-2)=-11y+17\]\[-10m+15-y+2=-11y+17\]\[-10m-y+17=-11y+17 \]\[-10m-y=-11y\]\[-10m=-10y\]\[m=y\] So what this tells you is that m and y must always be the same number for this equation to be true. As in the example I gave you (m=5,y=5) this was true.

OpenStudy (anonymous):

ok the first part before the equal sign was the problem, the second part was the answer. 11y+17. I missed 2 days so I have no clue about y and m's value. This is the first time I've heard of this.

OpenStudy (anonymous):

Simplifying is easy otherwise but this is the first time I've had a problem with m and y. I missed out on the instructions.

OpenStudy (anonymous):

Oh, if you were only given the first half of that equation, then I have no idea how you would get to the second half. Maybe you're missing something.

OpenStudy (anonymous):

No I have all the paperwork. The problem was to simplify -5(2m-3)-(y-2) That was all....

OpenStudy (anonymous):

Then they gave us the answer -11y+17

OpenStudy (anonymous):

Well you can't just drop a variable (m), so maybe it's a typo.

OpenStudy (anonymous):

ok that's what I was thinking...so if you simplify this problem what do you come up with?

OpenStudy (anonymous):

\[-5(2m-3)-(y-2) \]\[-10m+15-y+2\]\[-10m-y+17\]That's really as far as you can simplify it.

OpenStudy (anonymous):

If the m was supposed to be a y, it would simplify to the answer they gave.

OpenStudy (anonymous):

Thank you. That is the answer I got too. I'll talk to the teacher tomorrow about a possible type 0. Thanks again.